OpenStudy (anonymous):

Identify the curve by finding its Cartesian equation. r = tan(theta) sec(theta) .

OpenStudy (anonymous):

\[r=\tan\theta\sec\theta=\frac{\sin\theta}{\cos\theta}\sec\theta=\frac{\sin\theta}{\cos^2\theta}=\frac{\sin\theta}{1-\sin^2\theta}\] Recall that \(y=r\sin\theta\), so \[r=\frac{\frac{y}{r}}{1-\frac{y^2}{r^2}}\\ r=\frac{\frac{y}{r}}{1-\frac{y^2}{r^2}}\cdot\frac{r^2}{r^2}\\ r=\frac{ry}{r^2-y^2}\\ 1=\frac{y}{r^2-y^2}\] Also recall that \(r^2=x^2+y^2\): \[1=\frac{y}{x^2+y^2-y^2}~~\Rightarrow~~1=\frac{y}{x^2}~~\Rightarrow~~y=x^2\]